# Article Category

### v2

@@ 5.3.1 COMPUTATION
The distances to the feasible action sets can be computed equivalently if they are general zonotopes. For axis-aligned boxes, this can be computed explicitly. For explicitly available safe action zonotopes, the safe centre is defined as the centre of the zonotope. For induced safe action sets the safe centre is not readily available. We present two approaches to approximate it: orthogonal and zonotopic approximation. They are visualised in Figure˜4.
Figure 4: Approximation of the safe centre by expanding a contained zonotope (left) and by piercing the safe action set orthogonal to the boundary and taking the midpoint (right).
The zonotopic approach directly approximates the safe action set by inflating the generator lengths of a zonotope. The under-approximated zonotope $\mathcal{Z}_{\mathcal{A}_{s}}$ is the solution to
$\!\max_{c_{\mathcal{A}_{s}},l_{s}}$ $n]{\prod_{i=1}^{n}l_{s,i}}$ (23a)
subject to $\mathcal{Z}_{\mathcal{A}_{s}}=\left<c_{\mathcal{A}_{s}},G_{\mathcal{A}_{s}}\text{diag}(l_{s})\right>$ (23b)
@@
with $n$ uniformally sampled generator directions $G_{\mathcal{A}_{s}}$. Generally, the number of generators should be in the order of magnitude of the action dimension to provide a good approximation. However, $n$ should also not be too large, since we employ the geometric mean as a computationally cheaper proxy for the volume [12]. The issue arises from the uniformly sampled generator directions, which will likely lead to nearly parallel generators for $n\gg d$. However, computing the volume is only equivalent to the geometric mean for orthogonal generators, since the hypercube volume is the $n$th power of the geometric mean. The geometric mean favours spherical zonotopes for nearly parallel generators over elongated ones, which is not necessarily volume-maximising if the proper safe action set is elongated.
The orthogonal approximation does not directly approximate the safe action set. Instead, it pierces the safe action set orthogonal to the boundary and assumes the midpoint as the safe centre. The orthogonal starting point and direction is determined by Equation˜12, which yields $a_{s,\text{BP}}$ and $d_{\perp}=\frac{a_{s,\text{BP}}-a}{\left\lVert a_{s,\text{BP}}-a\right\rVert}$. Next, we reuse Equation˜22 as
@@
Then, the middle point is the safe centre
@@
Since Equation˜12 will only yield a different action for unsafe actions, the orthogonal approximation technique is restricted to those and actions that are already safe remain unmodified.
5.3.2 PROPERTIES
The generalised ray mask fulfils Property˜P1, since its codomain is the safe action set. To illuminate this fact, we remark that the function can be examined in one dimension, the direction along the ray, without loss of generality. The two limit points on the ray are $c_{\mathcal{A}_{s}}$, which maps to $g_{\mathcal{A}_{s},\text{RM}}(c_{\mathcal{A}_{s}})=c_{\mathcal{A}_{s}}\in\mathcal{A}_{s}$, and $c_{\mathcal{A}_{S}}+\lambda_{\mathcal{A}}d_{a}$, which maps to $g_{\mathcal{A}_{s},\text{RM}}(c_{\mathcal{A}_{S}}+\lambda_{\mathcal{A}}d_{a})=c_{\mathcal{A}_{s}}+\lambda_{\mathcal{A}_{s}}d_{a}\in\mathcal{A}_{s}$. Gradients to fulfil Property˜P2 are available from backpropagating through Equation˜20 and the implicit function theorem yields $\frac{\partial c_{\mathcal{A}_{s}}}{\partial a}$ $\frac{\partial\lambda_{\mathcal{A}}}{\partial a}$, and $\frac{\partial\lambda_{\mathcal{A}_{s}}}{\partial a}$.
Regarding smoothness, the ray mask safeguard is of class $C^{1}$ almost everywhere except for the safe set edges, as in boundary projection. The Jacobian of a ray mask is full-rank, adhering mostly to Property˜P3.
Lemma 3.
Let $\mathcal{A}_{s}$ be convex. Then, the Jacobian of any ray mask, as in Equation˜20, has full rank.
We obtain a proof by transforming the action space to a spherical coordinate system, which we provide in Section˜7.9. While the ray mask propagates gradient components in the mapping direction, they are still diminished for the linear mapping. This reduction is most obvious in the scenario, where the feasible and safe action set are spheres with coinciding centres and radii $r_{\mathcal{A}}>r_{\mathcal{A}_{s}}$, and the coordinate system is already spherical and centred. In this scenario, the Jacobian in Equation˜76 reduces to
$\frac{\partial a_{s}}{\partial a}=\begin{bmatrix}{\color[rgb]{0,0,0}\frac{r_{\mathcal{A}_{s}}}{r_{\mathcal{A}}}}&\textbf{0}\\ \textbf{0}&I\end{bmatrix}\,$ (26)
@@ 5.3.2 PROPERTIES
Contrary to the boundary projection safeguard, the ray mask applies to all actions, including safe actions. Moreover, the linear mapping distance decreases only linearly with the distance to the safe centre, as the partial derivative is constant in $\lambda_{a}$
@@
This means that safe actions far from the centre are also substantially altered, therefore Property˜P4 is not strongly adhered.
In regards to Property˜P5 and computational complexity, the actual application of the ray mask in Equation˜20 is negligible. However, computing the safe boundary Equation˜22 is a linear problem, since Equation˜22b is Equation˜7 or Equation˜8 depending on the availability of the safe action set. The safe centre is given by an explicit safe action set, but for induced ones, the approximations can also be costly. For a linearised transition distribution, the zonotopic approach is a conic problem, while the orthogonal approximation requires the solution of one quadratic problem.
5.3.3 MODIFICATIONS
We propose three possible modifications to the linear ray mask to improve its learning properties. First, we can increase the gradient in the mapping direction with the same regularisation term as in Equation˜15 to compensate for the diminished gradient and nudge towards safety.
Second, we can replace the Jacobian with an identity matrix for faster computation and unimpeded gradient propagation. An identity Jacobian retains the correct gradient directions if the reward-maximising action is safe. However, whether the reward-maximising action is safe is generally unknown and depends on the environment. For unsafe reward-maximising actions, the point of convergence would be the safe boundary point on the line $\overline{c_{\mathcal{A}_{s}}a_{r_{max}}}$, which is no longer optimal.
Finally, we propose the hyperbolic mapping function to limit the perturbation of safe actions. We visually compare both mappings in Figure˜5.
Figure 5: One-dimensional illustration of the linear ray mask safeguarding (RM) and the hyperbolic ray mask safeguarding (HRM). Arrows show how exemplary unsafe actions are mapped to the corresponding safeguarded actions. Shaded regions indicate the distribution of safe actions assuming an initially uniform distribution before safeguarding. The linear ray mask maps $\mathcal{A}$ linearly onto $\mathcal{A}_{s}$, whereas the hyperbolic ray mask maps it exponentially, such that unsafe actions are projected closely to the safe boundary and safe actions are perturbed minimally.
@@ 5.3.3 MODIFICATIONS
$\omega_{\text{tanh}}(\lambda_{a},\lambda_{\mathcal{A}_{s}},\lambda_{\mathcal{A}})=\frac{\tanh{\frac{\lambda_{a}}{\lambda_{\mathcal{A}_{s}}}}}{\tanh{\frac{\lambda_{\mathcal{A}}}{\lambda_{\mathcal{A}_{s}}}}}\,$ (28)
which maps unsafe actions close to the boundary and safe actions close to themselves, since $\omega_{\text{tanh}}(\lambda_{a}\approx\lambda_{\mathcal{A}_{s}})\approx\lambda_{\mathcal{A}_{s}}$ and $\omega_{\text{tanh}}(\lambda_{a}\ll\lambda_{\mathcal{A}_{s}})\approx\frac{\lambda_{a}}{\lambda_{\mathcal{A}_{s}}}$. It is a valid mapping as $w_{\text{tanh}}(\lambda_{a}=\lambda_{\mathcal{A}})=1$, $w_{\text{tanh}}(\lambda_{a}=0)=0$, and
@@ 6 NUMERICAL EXPERIMENTS
    Enabling our modified safeguards for analytic gradient-based reinforcement learning during training leads to similar or better policies than unsafe training, given the same number of environment interactions.
@@
6.1 SETUP
We conducted all experiments using ten different random seeds. Hyperparameters were tuned exclusively for non-safeguarded training and remained consistent throughout the safeguarded experiments [70]. We assessed the quality of the final policy by calculating the return (Return) achieved over a representative evaluation set. We tracked the number of steps until convergence
(# Steps), defined as reaching within 5% of the final policy. Further, we report the mean and a 95% confidence interval computed using bootstrapping for both the return and number of steps. Lastly, we report the number of runs that did not converge within the maximum allowed number of environment interactions (# Stuck). We excluded non-convergent runs from the return and step calculations to ensure clarity.
@@ 6.1.1 LEARNING ALGORITHMS
6.1.2 SAFEGUARDS
We evaluate the base versions of the boundary projection and ray mask as safeguards, where we approximate the safe centre using the zonotopic approach. We also assess all modifications to the safeguards individually, as well as the combination of regularisation and the hyperbolic ray mask. We did not investigate all combinations of modifications due to computational constraints.
6.1.3 ENVIRONMENTS
We study three model problems which are detailed in Section˜7.1. The first two are balancing tasks for a pendulum and a quadrotor, where we minimise the distance to an equilibrium position. The safety constraints comprise both action and state constraints that limit, for example, angles and angular velocities. To guarantee constraint satisfaction for all times, we use robust control invariant sets [46] as time-invariant safe state sets. The third environment features an energy management system for a battery and a heat pump aimed at minimising the electricity cost of a household while maintaining a comfortable indoor temperature. Both the state of charge of the battery and the room temperature have limits that must be enforced at all times. When considering the full action range, we achieve this by computing a safe state set that ensures the system can be steered back into the feasible set within one time step.
@@ 6.1.3 ENVIRONMENTS
6.2 EVALUATION OF LEARNING ALGORITHMS
In this subsection, we evaluate Hypothesis˜H1 by comparing the sampling-based reinforcement learning algorithms PPO and SAC with the analytic gradient-based algorithm SHAC in safeguarded training.
We first compared the learning algorithms in unsafe training to establish a baseline. We display the key metrics in Table˜2 and the learning curves in Section˜7.2. SHAC converged to the best policies in the pendulum and quadrotor tasks, where it was the only algorithm to balance the quadrotor consistently with minimal effort. However, it performed substantially worse on the energy system task. PPO performed best in energy systems but worst in the pendulum and quadrotor environments.
We attribute the performance degradation of SHAC in the energy system task to the high stochasticity of the environment. Environmental noise likely disrupts the computation of meaningful analytic gradients, and the smooth surrogate critic employed by SHAC may poorly approximate the true reward landscape. In contrast, PPO does not rely on a smooth reward approximation and benefits from the large number of simulation interactions available in the energy system task. Nevertheless, both PPO and SHAC had runs that failed to learn a meaningful policy in the energy system environment. This was also the case for one SHAC run in the pendulum environment.
@@ 6.2 EVALUATION OF LEARNING ALGORITHMS
Pendulum SHAC 12800 [10808, 15360] -8 [-8, -8] 1 / 10
SAC 8513 [6260, 10767] -14 [-15, -11] 0 / 10
PPO 81600 [81600, 81600] -596 [-1174, 300] 0 / 10
Quadrotor SHAC 20364 [11558, 28070] -157 [-169, -140] 0 / 10
SAC 80628 [60096, 109174] -1046 [-1855, 170] 0 / 10
PPO 80640 [42560, 116480] -1710 [-2128, -1267] 0 / 10
@@
SAC 674999 [554974, 781998] -5225 [-9424, 353] 0 / 10
PPO 748800 [645120, 861120] -2739 [-4598, -167] 2 / 10
After establishing the baselines in unsafe training, we proceed to the hypothesis, comparing safeguarded training. We show the key metrics in Table˜3 and the learning curves in Section˜7.3 and Section˜7.4. On the pendulum task, SAC initially reached near-optimal performance within the first evaluation but later diverged. Under uninformed safeguarding, SAC should benefit from its off-policy nature, but its reliance on the probability of the chosen action outweighs this effect. This issue was most noticeable in the critic loss, which was continuously divergent. The reported lower number of steps for SAC in the pendulum environment is an artefact of the inferior policy. On the quadrotor task, SAC learned ineffectively until the buffer reset roughly twice, at which point a jump in performance was consistently visible. The poor initial performance could result from uninformative earlier samples, although the underlying reason for the drastic performance increase is unclear. PPO benefited from the guided exploration in the pendulum and quadrotor environments, as safety is strongly tied to reward. In the energy system environment, ray masking for PPO significantly hinders learning. This is likely attributed to optimal actions often lying on the boundary of the safe action set in this task. The non-convergent runs of SHAC with boundary projection can be attributed to a total loss of gradient information as outlined in Equation˜14 since the action space in the pendulum environment is one-dimensional. Our observations support the hypothesis since the final policy and convergence speed of SHAC remained superior in the pendulum and quadrotor, while it narrowed the gap in the energy system.
@@
Pendulum BP SHAC 23360 [18560, 28800] -8 [-8, -8] 2 / 10
SAC 2504 [2504, 2504] -1083 [-1103, -1061] 0 / 10
PPO 80240 [78880, 82960] -10 [-10, -9] 0 / 10
RM SHAC 27392 [20992, 33280] -8 [-8, -8] 0 / 10
SAC 2504 [2504, 2504] -424 [-465, -384] 0 / 10
PPO 76160 [72080, 80240] -12 [-12, -11] 0 / 10
Quadrotor BP SHAC 45683 [16498, 74854] -333 [-394, -265] 0 / 10
SAC 110176 [90131, 123196] -338 [-368, -308] 0 / 10
PPO 118720 [89600, 152320] -402 [-453, -350] 0 / 10
RM SHAC 67148 [31909, 99636] -251 [-307, -197] 0 / 10
SAC 80128 [59081, 107171] -377 [-415, -337] 0 / 10
PPO 127680 [107520, 156800] -379 [-419, -330] 0 / 10
Energy System BP SHAC 366960 [213807, 509553] -89167 [-111791, -62775] 0 / 10
SAC 491000 [290000, 685050] -150179 [-252908, -33560] 0 / 10
PPO 661577 [457426, 905307] -2293 [-3421, -906] 3 / 10
RM SHAC 709280 [579920, 840433] -8793 [-12685, -3679] 0 / 10
SAC 355999 [187974, 503073] -1843 [-2396, -1279] 0 / 10
@@
6.3 EVALUATION OF SAFEGUARDS
Next, we evaluate Hypothesis˜H2 by comparing the various safeguards introduced in Section˜5 to unsafe training on SHAC. Figure˜6 shows the aggregated learning curves; we report the number of non-convergent runs in Table˜4. Compared to unsafe training, we expect performance to decline for the unaltered safeguards when the optimal action is safe, as is the case for most actions in balancing scenarios. The impact should be more severe for the unaltered boundary projection than for the ray mask, attributed to the lack of gradient propagation in the mapping direction.
Figure 6: Comparison of SHAC in unsafe and safeguarded training via boundary projection (left) and ray mask (right). The modifications are indicated by a different line style. At least one of our safeguards achieves comparable performance to unsafeguarded training.

### v3

@@ 5.C.1 COMPUTATION
The distance to the feasible action set $\lambda_{\mathcal{A}}$ can be computed equivalently if it is a zonotope. For an axis-aligned box, the computation is possible in closed-form [73]. For a specified safe action zonotope, the safe centre is defined as the centre of the zonotope. The safe centre is not readily available for a derived safe action set, as in Equation˜2. We present two approaches to approximate it: orthogonal and zonotopic approximation, which are visualised in Figure˜4.
[…]
Figure 4: Zonotopic approximation of the safe centre $c_{\mathcal{A}_{s}}$ by expanding a contained zonotope $\mathcal{Z}_{\mathcal{A}_{s}}$ (left) and orthogonal approximation by piercing the safe action set $\mathcal{A}_{s}$ orthogonal to the boundary and taking the midpoint between both boundary points $a_{s,\text{BP}}$ and $a_{s,\text{BP}}+\lambda_{\perp}d_{\perp}$ (right).
The zonotopic approach directly approximates the safe action set by maximising the generator lengths of a zonotope while maintaining containment. The under-approximated zonotope $\mathcal{Z}_{\mathcal{A}_{s}}$ is the solution to
$\!\max_{c_{\mathcal{A}_{s}},l_{s}}$ $\prod_{i=1}^{n}l_{s,i}$ (26a)
subject to $\mathcal{Z}_{\mathcal{A}_{s}}=\left<c_{\mathcal{A}_{s}},G_{\mathcal{A}_{s}}(l_{s})_{D}\right$> (26b)
@@
with $n$ generator directions $G_{\mathcal{A}_{s}}$ sampled uniformly from a $d$-dimensional sphere $\mathbb{S}^{d}$ and where we denote the diagonalisation of a vector by the subscript $D$. Generally, the number of generators should be in the order of magnitude of the action dimension to provide a good approximation. However, $n$ should also not be too large, since we employ the volume computation of a box as a computationally cheaper proxy for the volume of a zonotope [12], which assumes orthogonal generators. This assumption is violated for $n>d$. Therefore, the objective Equation˜26a favours spherical zonotopes over elongated ones, which is not necessarily volume-maximising if the proper safe action set is elongated.
The orthogonal approximation computes the required safe centre $c_{\mathcal{A}_{s}}$ and distances $\lambda_{a}$, $\lambda_{\mathcal{A}}$, and $\lambda_{\mathcal{A}_{s}}$ without the expensive approximation of the safe action set. Instead, it pierces the safe action set orthogonal to the boundary and assumes the midpoint between the entry and exit point as the safe centre. The orthogonal starting point and direction is determined by Equation˜12, which yields $a_{s,\text{BP}}$ and $d_{\perp}=\frac{a_{s,\text{BP}}-a}{\left\lVert a_{s,\text{BP}}-a\right\rVert_{2}}$. Next, we reuse Equation˜25 as
@@
We utilise the midpoint between $a_{s,\text{BP}}$ and $a_{s,\text{BP}}+\lambda_{\perp}d_{\perp}$ as the safe centre
@@
Since Equation˜12 will only yield a different action for unsafe actions, the orthogonal approximation technique is restricted to those actions.
5.C.2 PROPERTIES
The generalised ray mask satisfies Property˜P1, since its codomain is the safe action set. To illuminate this fact, we remark that Equation˜21 can be examined in one dimension – the direction along the ray $d_{a}$ – without loss of generality. The action along this ray is bounded between $c_{\mathcal{A}_{s}}$, which maps to $g_{\text{RM}}(c_{\mathcal{A}_{s}})=c_{\mathcal{A}_{s}}\in\mathcal{A}_{s}$, and $c_{\mathcal{A}_{S}}+\lambda_{\mathcal{A}}d_{a}$, which maps to $g_{\text{RM}}(c_{\mathcal{A}_{S}}+\lambda_{\mathcal{A}}d_{a})=c_{\mathcal{A}_{s}}+\lambda_{\mathcal{A}_{s}}d_{a}\in\mathcal{A}_{s}$. Gradients to obtain Property˜P2 are available from backpropagating through Equation˜21.
Regarding smoothness, the ray mask safeguard is of class $C^{1}$ almost everywhere, except for the safe set edges and for the $\epsilon$-sphere around the safe action set centre $\mathcal{A}_{\epsilon}=\{a\in\mathcal{A}_{s}\mid\left\lVert a-c_{\mathcal{A}_{s}}\right\rVert_{2}=\epsilon\}$. As in boundary projection, we employ any element from the Clarke subdifferentiable [70] at these edges. The Jacobian of a ray mask has full rank, wherever $g_{\text{RM}}$ is differentiable. Consequently, a ray mask satisfies Property˜P3 almost everywhere.
Theorem 3.
Let $\mathcal{A}_{s}$ be convex, $g_{\text{RM}}$ differentiable and $\left\lVert a-c_{\mathcal{A}_{s}}\right\rVert_{2}>\epsilon$. Then, the Jacobian of any ray mask as in Equation˜21 has full rank.
We present the proof in Appendix˜C. While the ray mask propagates gradients in the mapping direction, they are still diminished for the linear mapping. This reduction is particularly obvious for the linear mapping function in the scenario where the feasible and safe action set are spheres with coinciding centres and radii $r_{\mathcal{A}}>r_{\mathcal{A}_{s}}$, and the coordinate system is already spherical and centred. In this scenario, the Jacobian in Equation˜77 reduces to
$\frac{\partial a_{s}}{\partial a}=\begin{bmatrix}\frac{r_{\mathcal{A}_{s}}}{r_{\mathcal{A}}}&\textbf{0}\\ \textbf{0}&I\end{bmatrix}\,$, (29)
@@ 5.C.2 PROPERTIES
Contrary to the boundary projection safeguard, the ray mask applies to all actions, including safe actions. Moreover, the linear mapping distance decreases only linearly with the distance to the safe centre, as the partial derivative is constant in $\lambda_{a}$:
@@
This means that safe actions far from the centre are also substantially altered, therefore Property˜P4 is not firmly adhered to.
In regard to Property˜P5 and computational complexity, the actual application of the ray mask in Equation˜21 is negligible, as it is a closed-form expression. However, computing the safe boundary Equation˜25 is a linear program, since Equation˜25b has to be considered through Equation˜7 or Equation˜8, depending on the availability of the safe action set. A specified safe action set provides the safe centre. However, for derived safe action sets, as in Equation˜2, the approximations can be costly. For linearised dynamics, the zonotopic approach is a conic program, while the orthogonal approximation requires the solution of one quadratic program.
5.C.3 MODIFICATIONS
We propose three possible modifications to the linear ray mask to improve its learning properties. First, we can increase the gradient in the mapping direction with the same regularisation term as in Equation˜16 to compensate for the diminished gradient and nudge towards safety.
Second, we can replace the Jacobian with an identity matrix for faster computation and unimpeded gradient propagation, which we denote passthrough. This modification retains the correct gradient directions if the reward-maximising action is safe. However, whether the reward-maximising action is safe is generally unknown and depends on the environment. For unsafe reward-maximising actions, the point of convergence of the policy optimisation would be the safe boundary point on the line $\overline{c_{\mathcal{A}_{s}}a_{r_{max}}}$, which is no longer optimal.
Finally, we propose a hyperbolic mapping function to limit the perturbation of safe actions. We visually compare both mappings in Figure˜5.
Figure 5: One-dimensional illustration of the linear ray mask (RM) and the hyperbolic ray mask (HRM). Arrows show mappings for exemplary unsafe actions to the corresponding safeguarded actions. Shaded regions indicate the distribution of safe actions. The linear ray mask maps $\mathcal{A}$ linearly onto $\mathcal{A}_{s}$, whereas the hyperbolic ray mask maps it exponentially, such that unsafe actions are projected closely to the safe boundary and safe actions are perturbed minimally.
@@ 5.C.3 MODIFICATIONS
$\omega_{\text{tanh}}(\lambda_{a},\lambda_{\mathcal{A}_{s}},\lambda_{\mathcal{A}})=\frac{\tanh{\frac{\lambda_{a}}{\lambda_{\mathcal{A}_{s}}}}}{\tanh{\frac{\lambda_{\mathcal{A}}}{\lambda_{\mathcal{A}_{s}}}}}\,$, (31)
which maps unsafe actions close to the boundary and safe actions close to themselves, since $\omega_{\text{tanh}}(\lambda_{a}>\lambda_{\mathcal{A}_{s}})\approx\lambda_{\mathcal{A}_{s}}$ and $\omega_{\text{tanh}}(\lambda_{a}<\lambda_{\mathcal{A}_{s}})\approx\frac{\lambda_{a}}{\lambda_{\mathcal{A}_{s}}}$. It is a valid mapping, as defined in Equation˜22 and Equation˜23, as $w_{\text{tanh}}(\lambda_{a}=\lambda_{\mathcal{A}})=1$, $w_{\text{tanh}}(\lambda_{a}=0)=0$, and
@@ 6 NUMERICAL EXPERIMENTS
    Enabling our modified safeguards for analytic gradient-based reinforcement learning during training leads to similar or higher return policies than unsafe training, given the same number of environment interactions.
@@
6.A SETUP
We conducted all experiments using ten different random seeds. Hyperparameters were tuned exclusively for non-safeguarded training and carried over unchanged to the safeguarded experiments [74]. We assessed the quality of the final policy by calculating the return (Return) achieved over a representative evaluation set. We tracked the number of steps until convergence (# Steps), defined as reaching within 5% of the return of the final policy. Further, we report the mean and a 95% confidence interval computed using bootstrapping for both the return and number of steps. Lastly, we report the number of runs that did not converge within the maximum allowed number of environment interactions (# Stuck). We excluded non-convergent runs from the return and step calculations to ensure clarity.
@@ 6.A.1 LEARNING ALGORITHMS
6.A.2 SAFEGUARDS
We evaluate the base versions of the boundary projection and ray mask as safeguards, where we approximate the safe centre using the zonotopic approach. We also assess all modifications to the safeguards individually, as well as the combination of regularisation and the hyperbolic ray mask.
6.A.3 ENVIRONMENTS
We study three environments, which are detailed in Appendix˜A. The first two are balancing tasks for a pendulum and a quadrotor, where we minimise the distance to an equilibrium position. The safety constraints comprise both action and state constraints that limit, for example, angles and angular velocities. To guarantee constraint satisfaction at all times, we use robust control invariant sets [46] as time-invariant safe state sets. The third environment features an energy management system for a battery and a heat pump aimed at minimising the electricity cost of a household while maintaining a comfortable indoor temperature. Both the state of charge of the battery and the room temperature have limits that must be enforced at all times. When considering the full action range, we achieve this by computing a safe state set that ensures that the system can be steered back into the feasible set within one time step.
@@ 6.A.3 ENVIRONMENTS
6.B EVALUATION OF LEARNING ALGORITHMS
In this subsection, we evaluate Hypothesis˜H1 by comparing the sampling-based reinforcement learning algorithms PPO and SAC with the analytic gradient-based algorithm SHAC.
We first compared the learning algorithms in unsafe training to establish a baseline. The key metrics are listed in Table˜2 and the learning curves in Appendix˜D. SHAC converged to the best policies in the pendulum and quadrotor tasks, where it was the only algorithm to balance the quadrotor consistently with minimal effort. However, it performed substantially worse on the energy system task. PPO performed best in energy systems but worst in the pendulum and quadrotor environments.
We attribute the performance degradation of SHAC in the energy system task to the high degree of noise of the environment. Environmental noise likely disrupts the computation of meaningful analytic gradients, and the smooth surrogate critic employed by SHAC may poorly approximate the true reward landscape. In contrast, PPO does not rely on a smooth reward approximation and benefits from the large number of simulation interactions available in the energy system task. Nevertheless, both PPO and SHAC had runs that failed to learn a meaningful policy in the energy system environment. This was also the case for one SHAC run in the pendulum environment.
@@ 6.B EVALUATION OF LEARNING ALGORITHMS
Pendulum SHAC 12800 [010808, 015360] -8 [00000-8, 0000-8] 1 / 10
SAC 8513 [006260, 010767] -14 [0000-15, 000-11] 0 / 10
PPO 81600 [081600, 081600] -596 [00-1174, -00300] 0 / 10
Quadrotor SHAC 20364 [011558, 028070] -157 [000-169, 00-140] 0 / 10
SAC 80628 [060096, 109174] -1046 [00-1855, -00170] 0 / 10
PPO 80640 [042560, 116480] -1710 [00-2128, 0-1267] 0 / 10
@@
SAC 674999 [554974, 781998] -5225 [00-9424, -00353] 0 / 10
PPO 748800 [645120, 861120] -2739 [00-4598, 00-167] 2 / 10
After establishing the baselines in unsafe training, we proceed to testing the hypothesis by comparing the learning algorithms in safeguarded training. We show the key metrics in Table˜3 and the learning curves in Appendix˜E and Appendix˜F. We obtained results similar to unsafe training, as SHAC converged to the best policies in the balancing tasks, whereas the sampling-based methods outperformed SHAC on the energy system. However, SHAC performed substantially better in safeguarded training in the energy system task.
SAC in safeguarded training showed volatile behaviour in balancing scenarios. For example, in the pendulum task, SAC initially reached near-optimal performance within the first evaluation but later diverged. In the quadrotor environment, SAC learned ineffectively until the buffer reset roughly twice, at which point a jump in performance was consistently visible. Under uninformed safeguarding, SAC should benefit from its off-policy nature, but its reliance on the probability of the chosen action outweighs this effect. This issue was most noticeable in the pendulum task, where the critic loss was continuously divergent. The poor initial performance in the quadrotor task could result from uninformative earlier samples, although the underlying reason for the drastic performance increase is unclear.
PPO mostly benefitted from safeguarded training, especially in the balancing environments. There, safety is strongly tied to reward, enabling safeguarding to guide the exploration. In the energy system environment, boundary projection had a similar effect, however, ray masking significantly hindered learning. We attribute this to the diminished learning rate in the ray direction.
The performance of SHAC with unaltered safeguards was mostly similar to unsafe training. Notable exceptions were the impaired learning in the quadrotor environment and the improved performance on the energy system task. The optimal action is mostly safe in balancing scenarios, such that the lack of gradient propagation in the mapping direction hurts learning. Due to the simplicity of the pendulum environment, the convergent runs showed barely any degradation compared to unsafe training. However, the increase in non-convergent runs on the pendulum with boundary projection is caused by a total loss of gradient information as outlined in Equation˜14, since the action space in the pendulum environment is one-dimensional. In the more complex quadrotor task, the agents were learning very slowly or completely stalled for several individual runs. In contrast, the increase in return in the energy system environment could be due to the diminished gradients, which stabilise learning there.
Our observations support Hypothesis˜H1 since the final policy and convergence speed of SHAC remained superior in the pendulum and quadrotor, while it narrowed the gap in the energy system.
@@
Pendulum BP SHAC 23360 [018560, 028800] -8 [00000-8, 00000-8] 2 / 10
SAC 2504 [002504, 002504] -1083 [00-1103, 00-1061] 0 / 10
PPO 80240 [078880, 082960] -10 [0000-10, 00000-9] 0 / 10
RM SHAC 27392 [020992, 033280] -8 [00000-8, 00000-8] 0 / 10
SAC 2504 [002504, 002504] -424 [000-465, 000-384] 0 / 10
PPO 76160 [072080, 080240] -12 [0000-12, 0000-11] 0 / 10
Quadrotor BP SHAC 45683 [016498, 074854] -333 [000-394, 000-265] 0 / 10
SAC 110176 [090131, 123196] -338 [000-368, 000-308] 0 / 10
PPO 118720 [089600, 152320] -402 [000-453, 000-350] 0 / 10
RM SHAC 67148 [031909,0 99636] -251 [000-307, 000-197] 0 / 10
SAC 80128 [059081, 107171] -377 [000-415, 000-337] 0 / 10
PPO 127680 [107520, 156800] -379 [000-419, 000-330] 0 / 10
Energy System BP SHAC 366960 [213807, 509553] -89167 [-111791, 0-62775] 0 / 10
SAC 491000 [290000, 685050] -150179 [-252908, 0-33560] 0 / 10
PPO 661577 [457426, 905307] -2293 [00-3421, 000-906] 3 / 10
RM SHAC 709280 [579920, 840433] -8793 [0-12685, 00-3679] 0 / 10
SAC 355999 [187974, 503073] -1843 [00-2396, 00-1279] 0 / 10
@@
6.C EVALUATION OF SAFEGUARDS
Next, we evaluate Hypothesis˜H2 by comparing the safeguards introduced in Section˜5 to unsafe training on SHAC. Figure˜6 shows the aggregated learning curves; we report the number of non-convergent runs in Table˜4.
For the unaltered safeguards, we observed a performance decline when the optimal action is safe compared to unsafe training. The impact was more severe for the unaltered boundary projection than for the ray mask, attributed to the lack of gradient propagation in the mapping direction. To this end, regularisation mostly improved the performance for both boundary projection and ray masking. Ray masking with a passthrough gradient and the hyperbolic ray mask had mixed results.
Figure 6: Comparison of SHAC under unsafe training and safeguarded training using boundary projection (left) and ray mask (right). Different line styles indicate the safeguard variants. In the pendulum environment, both base safeguards achieve comparable performance; in the quadrotor environment, regularised boundary projection and the regularised hyperbolic ray mask perform similarly; and in the energy system environment, all ray masks yield superior performance.